OpenStudy (ammarah):

Solve the equation: 25x^2 -24x-9=-7x^2 +12x-18

sam (.sam.):

What's the problem? Any ideas?

OpenStudy (ammarah):

um i made it equal to zero but idk how to factor it because i got 32x^2 -36x+9=0 i dont know what to do now....

sam (.sam.):

This one's a bit difficult to factorise, but think about two numbers that multiplied to get 32. Also think about another two numbers multiplied to get 9.

OpenStudy (ammarah):

i did think but i dont know what it is can u please show me step by step?

sam (.sam.):

For 32, you got (8 times 4) you get 32. For 9, try (3 times 3) you get 9. Now we have 8 and 4, 3 and 3 (8x ? 3)(4x ? 3)=0 The missing slot there is where you either put in + or - depends on the middle value, which we have there is -36x. Distributing that, So 8x times 3 is 24x 3 times 4x is 12x To get -36x, both 24x and 12x must be negative, so which means (8x - 3)(4x - 3)=0

OpenStudy (ammarah):

okay i have another question.. it says find the zeroes of the function by rewriting the function in intercept form. y=3x^2 +2x

sam (.sam.):

Finding zeroes means letting y=0 3x^2 +2x=0 x(3x+2)=0 x=0, x=-2/3

OpenStudy (ammarah):

wait no whyd u delete it? i need the work on that one

OpenStudy (anonymous):

I copied the problem incorrectly. I think Sam is correct.

OpenStudy (anonymous):

I'll rework the problem.

OpenStudy (ammarah):

okay i have another question... find wheter the graph has a maximum or minimum value find the value.. y= (x-6)^2+3

OpenStudy (anonymous):

32x^2-36x+9=0 (8x-3)(4x-3)=0 x= 3/8 or x= 3/4

OpenStudy (ammarah):

no its graphing read the question

OpenStudy (anonymous):

y= (x-6)^2+3 When I graphed, your have a minimum (6,3) You can find a free graphing calc at http://www.meta-calculator.com/online/

OpenStudy (anonymous):

25x^2-24x-9=-7x^2+12x-18 (25x^2+7x^2)+(-24x-12x)+(-9+18)=0 32x^2-36x+9=0 x^2-36/32x+9/36=0 x^2-9/8x+1/4=0 x^2-81/256+81/256-9/8x+1/4=0 x^2-9/8x+81/256+(1/4-81/256)=0 (x-9/16)^2=81/256-1/4 x-9/16=sqrt(81/256-1/4) x=sqrt(81/256-1/4)+9/16

sam (.sam.):

For y= (x-6)^2+3, y= (x-a)^2+b Vertex of the graph is (a,b), so the vertex is (6,3) --------------------------------------- When you expand this (x-6)^2+3 , you should get positive coefficient for x^2, that means the graph should be a minimum graph/ curve upwards.  ---------------------------------------- You could also find out where the y-intercept is by setting 'x' = 0, the y-intercept should be 39. y=(x-6)^2+3 y=(-6)^2+3 y=36+3 y=39